OpenStudy (unklerhaukus):

For the following DE, check that \(y_1\) is a solution of the corresponding homogeneous equation and then calculate the general solution by the reduction of order method. \[y^{\prime\prime} -4y^\prime +3y=3x-4;\qquad y_1 =e^x\]

OpenStudy (unklerhaukus):

\[y^{\prime\prime} -4y^\prime +3y=3x-4;\qquad y_1 =e^x\] \[\qquad\qquad\qquad\qquad\qquad\qquad y_1^\prime =e^x\]\[\qquad\qquad\qquad\qquad\qquad\qquad y_1^{\prime\prime} =e^x\] \[y_1^{\prime\prime} -4y_1^\prime +3y_1=0\]\[\left(e^x\right) -4\left(e^x\right)+3\left(e^x\right) =0\]\[\left(1-4+3\right)e^x=0\] \[\therefore\qquad y_1= e^x\qquad\text{is a solution to the homogenous equation}\]

OpenStudy (unklerhaukus):

\[y=y_p+y_c\]\[y_p=y_1+y_2\]\[y_2=uy_1\]

OpenStudy (unklerhaukus):

\[y_2=uy_1=ue^x\]\[y_2^\prime=u^\prime e^x+ue^x=(u^\prime+u)e^x\]\[y_2^{\prime\prime}=(u^{\prime\prime}+u^\prime)e^x+(u^\prime+u)e^x\]\[y_2^{\prime\prime}=(u^{\prime\prime}+2u^\prime+u)e^x\] \[y_2^{\prime\prime} -4y_2^\prime +3y_2=0\]\[\left((u^{\prime\prime}+2u^\prime+u)e^x\right) -4\left((u^\prime+u)e^x\right) +3\left(ue^x\right)=0\]\[\left(u^{\prime\prime}+2u^\prime+u -4u^\prime-4u +3u\right)e^x=0\]\[\left(u^{\prime\prime}-2u^\prime\right)e^x=0\]\[u^{\prime\prime}-2u^\prime=0\]

OpenStudy (unklerhaukus):

\[\text{let } p=u^\prime\]\[p^\prime=u^{\prime\prime}\] \[p^\prime-2p=0\]

OpenStudy (unklerhaukus):

\[\mu(x)=e^{\int-2\text dx}=e^{-2x}\] \[\left(pe^{-2x}\right)^\prime=0\]\[pe^{-2x}=\int 0\cdot\text dx\]\[pe^{-2x}=1\]\[p=e^{2x}\]\[u^{\prime}=e^{2x}\]\[u=\frac {e^{2x}}2\]

OpenStudy (unklerhaukus):

\[y_2=e^{2x}y_1\] \[y_p=ae^x+be^xe^{2x}\]\[=ae^x+be^{3x}\]

OpenStudy (unklerhaukus):

how should i find \(y_c\)

OpenStudy (anonymous):

Laplace transform is the best way to solve it. follow me.

OpenStudy (unklerhaukus):

the question is asking for reduction of order method.

OpenStudy (anonymous):

ok I will give you a beautiful solution to reduction. follow me.

OpenStudy (experimentx):

@mahmit2012 doesn't Laplace transform requite initial values??

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):